OpenStudy (anonymous):

Will give medal if you help!!!:D and will fan if i havent already!!!:D /) /) ( ^_^) C (") (")

OpenStudy (anonymous):

question?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

What is the solution to the system of equations represented by these two lines?

OpenStudy (anonymous):

do you have some sort of picture with the problem?

OpenStudy (anonymous):

A. (2, 3) B. (0, 4) C. (3, 2) D. (3, 0)

OpenStudy (anonymous):

c

OpenStudy (anonymous):

ok can i ask 4 more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations?

OpenStudy (anonymous):

dont worry hold on im posting up the other parts

OpenStudy (anonymous):

y=3/2x y=-1/2x+4

OpenStudy (anonymous):

this one has me stuck

OpenStudy (anonymous):

its ok i got it did a graph cheat but i can only do it a couple of times ;)

OpenStudy (anonymous):

ok! you have 3 more, correct?

OpenStudy (anonymous):

xD if i could i would fan myself

OpenStudy (anonymous):

and yes

OpenStudy (anonymous):

now this on might take a bit to get it up ok

OpenStudy (anonymous):

Ben graphed a system of equations and stated that the solution is (4, 4). Check his solution algebraically. Which statement is true?

OpenStudy (anonymous):

The graph of y=-1/2x+5 is incorrect.

OpenStudy (anonymous):

thats A.

OpenStudy (anonymous):

B. The graph of y=1/2x+2 is incorrect.

OpenStudy (anonymous):

C. (2, 4) is the solution to the system of equations. D. (4, 2) is the solution to the system of equations.

OpenStudy (anonymous):

C i believe

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What is the best approximation of the solution to the system to the nearest integer values? A. (1, 3) B. (3, 2) C. (2, 3) D. (3, 4)

OpenStudy (anonymous):

What is the best approximation of the solution to the system to the nearest integer values?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

there was no D.

OpenStudy (anonymous):

lol i passed it anyway :P if i were to of missed one more i wouldnt of but i did so im happy ^_^

OpenStudy (anonymous):

are you taking plato also? lol

OpenStudy (anonymous):

idek what that is xD

OpenStudy (anonymous):

the online school thing

OpenStudy (anonymous):

yeah, i like scooby-Doo